# Deep Neural Networks and Brain Alignment: Brain Encoding and Decoding (Survey)

### v2

@@ 7.6 Auditory Decoding
With the recent advancements of self-supervised speech models and generative AI models, recent studies have largely targeted reconstructing speech/music from brain recordings (Défossez et al., 2023; Denk et al., 2023; Senda et al., 2024). As shown in Figure 21Défossez et al. (2023) proposed a CLIP-MEG pipeline to align MEG activity onto pretrained speech embeddings and generate speech from a stream of MEG signals. Unlike other methods which are experimented with on narrative speech, Denk et al. (2023) introduce a method for reconstructing music from fMRI brain activity, as shown in Figure 22. Specifically, they proposed a Brain2Music pipeline where the first step involves using fMRI data to predict MuLanmusic embeddings (Huang et al., 2022), which are then passed to MusicLM (Agostinelli et al., 2023), is conditioned to generate the music reconstruction, resembling the original music stimulus.
Figure 21: CLIP-MEG pipeline to align MEG activity onto pretrained speech embeddings. The Figure is adapted from Défossez et al. (2023).
Figure 22: Brain2Music decoding pipeline adapted from (Denk et al., 2023)The Figure is adapted from Denk et al. (2023).
@@ 7.7.1 Image Reconstruction
Before the success of recent generative AI models, researchers have used deep-learning models and algorithms, including generative adversarial networks (GANs) and self-supervised learning models trained on a large number of naturalistic images (Du et al., 2020; Beliy et al., 2019; Fang et al., 2020; Gaziv et al., 2022; Lin et al., 2022b). For instance, Beliy et al. (2019) designed a separable autoencoder that enables self-supervised learning in fMRI and images to increase training data. Mind Reader (Lin et al., 2022a) encoded fMRI signals into a pre-aligned vision-language latent space and used StyleGAN2 (Karras et al., 2020) for image generation. These methods generate more plausible and semantically meaningful images. Several other studies focused on reconstructing personal imagined experiences (Berezutskaya et al., 2020) or application-based decoding like using brain activity scanned during a picture-based mechanical engineering task to predict individuals’ physics/engineering exam results (Cetron et al., 2019) and reflecting whether current thoughts are detailed, correspond to the past or future, are verbal or in images (Smallwood & Schooler, 2015).
@@
With the recent success of CLIP and Diffusion models, deep generative models have been gaining attention to generate high-resolution images with high semantic fidelity (Takagi & Nishimoto, 2023; Chen et al., 2023; Scotti et al., 2024; Benchetrit et al., 2023; Song et al., 2023).  (Takagi & Nishimoto, 2023) proposed a method for image reconstruction from fMRI using Stable Diffusion (Rombach et al., 2022), as shown in Figure 23(left). Their approach involves decoding brain activities to text descriptions and converting them to natural images using Stable Diffusion. Based on a similar philosophy, using a Stable Diffusion model as a generative prior and the pretrained fMRI features as conditions, (Chen et al., 2023) reconstructed high-fidelity images with high semantic correspondence to the groundtruth stimuli, as shown in Figure 23(right).  (Scotti et al., 2024) proposed a MindEye that can map fMRI brain activity to any high dimensional multimodal latent space, like CLIP image space, enabling image reconstruction using generative models that accept embeddings from this latent space. Different from previous studies, BrainCLIP framework was introduced by (Liu et al., 2023) to align fMRI patterns with different modalities (especially from visual and textual modalities) through cross-modal contrastive loss. All these studies have been limited to 2D visual representations. A recent work (Gao et al., 2023) aims to extend the scope of fMRI decoding to 3D representations. Specifically, Gao et al. (2023) introduce Recon3DMind, a groundbreaking task focused on reconstructing 3D visuals from fMRI signals.
@@
Lastly, recent image reconstruction studies have focused on other non-invasive brain recordings such as MEG and EEG rather than fMRI signals.  (Benchetrit et al., 2023) proposed a CLIP-MEG pipeline to align MEG activity onto pretrained visual embeddings and generate images from a stream of MEG signals. Similarly, (Song et al., 2023) proposed a CLIP-EEG pipeline to align these two modalities (image and EEG encoders to extract features from paired image stimuli and EEG responses) by constraining their similarity.
@@
Figure 23: Image reconstruction from fMRI using Stable Diffusion. Left: (Takagi & Nishimoto, 2023), Right: (Chen et al., 2024)The Left Figure is adapted from Takagi & Nishimoto (2023) and the Right Figure is adapted from Chen et al. (2024).
@@ 7.7.2 Video Reconstruction
Unlike static natural images, human visual cortex can process a continuous, diverse flow of scenes, motions, and objects. To recover dynamic visual experience, the challenge lies in the nature of fMRI, which measures blood oxygenation level dependent (BOLD) signals and captures snapshots of brain activity every few seconds. Similar to image reconstruction works, Chen et al. (2024) present MinD-Video, a two-module pipeline (i.e. CLIP module followed by latent stable diffusion) designed to bridge the gap between image and video brain decoding.
@@ 8 Conclusion, Limitations, and Future Trends
AI engineering. The recent brain encoding studies most immediately fit in with the neuro-AI research direction that specifically investigates the relationship between representations in the brain and representations learned by powerful neural network models. This direction has gained recent traction, especially in the domain of language, vision, and speech processing, thanks to advancements in language models (Schrimpf et al., 2021; Goldstein et al., 2022), vision models (Schrimpf et al., 2020) and speech models (Tuckute et al., 2023; Oota et al., 2023c). Furthermore, several recent works most immediately contribute to this line of research by understanding the reasons for the observed similarity in more depth (Merlin & Toneva, 2022; Oota et al., 2024b; Kauf et al., 2024b; Sarch et al., 2024; Oota et al., 2024a). Overall, these studies provide valuable insights for selecting features, enhancing transfer learning, and aiding in the creation of AI architectures that are cognitively plausible.
@@ 8.1 Future Trends
Bridging the Gap: Enhancing Deep Neural Network Models for Deeper Insights into Auditory, Language and Visual Processing While significant progress has been made in understanding text-based models, understanding the similarity in information processing between visual, speech, and multimodal models versus natural brain systems remains an open area. For instance, Oota et al. (2024a) demonstrate that speech-based language models lack brain relevant semantics in language regions. Therefore, enhancing speech-based language models to align more closely with text-based models could provide valuable insights into language and auditory processing, given that speech is the most ancient form of human language. This suggests a promising direction for future research, aiming to bridge the gap between artificial intelligence models and the complex, multifaceted processes of human cognition.
@@
Advancing Multimodal Decoding: The Next Leap in Deep Learning Accuracy Decoding actual multimodal stimuli has become increasingly feasible due to recent advancements in deep learning models dedicated to generation tasks (Rombach et al., 2022; Singer et al., 2022). However, there is still a significant need for further research to enhance the accuracy of these models. This involves not only refining the algorithms and architectures used but also improving the quality and diversity of the datasets on which these models are trained. Advancements in computational power, algorithmic efficiency, and innovative training methodologies are critical for pushing the boundaries of what is possible in multimodal decoding, aiming to achieve more precise, reliable, and nuanced interpretations of complex stimuli.
@@
Mapping the Mind: The Effects of Brain Damage on Cognitive Capabilities We need a deeper understanding of the degree to which damage to different regions of the human brain could lead to the degradation of selective cognitive skills. This exploration requires detailed mapping of cognitive functions to specific brain areas, taking into account the brain’s complex network of connections. Studies should investigate not only the immediate effects of brain damage on cognitive skills but also the brain’s capacity for reorganization and compensation over time. Ultimately, the goal is to translate these research findings into practical applications, such as more effective cognitive rehabilitation techniques and assistive technologies to improve the quality of life for individuals with brain injuries.
@@
Towards Human-Like Understanding in ANNs: Integrating Self-Supervised Learning and Brain-Inspired Architectures How can we train artificial neural networks in novel self-supervised ways such that they compose word meanings or comprehend images and speech like a human brain? Can we model the hierarchical and modular organization of the brain in neural network architectures? This involves creating networks that reflect the brain’s organization, from low-level feature detection to high-level semantic processing, allowing for the integration of information across different modalities. Moreover, how might we integrate dynamic learning strategies, such as curriculum learning, which progressively introduces more intricate tasks to the model? This method emulates how humans naturally progress from understanding straightforward to more complex ideas over time.
@@
Bridging the Language Gap in Brain-NLP Research: The Need for Multilingual Exploration An important part of brain-NLP research relies on brain recordings collected from individuals who speak English as their primary language. Additionally, these studies utilize experimental stimuli that are presented in the English language. As a result, all current neuro-AI studies predominantly leverages language models and neural models that have been trained extensively on English text data and brain responses elicited by text or speech in English. However, it is essential to acknowledge the potential variability in our study outcomes when extrapolated to languages other than English. The intricate interplay between language-specific nuances and neural responses may introduce distinctions in the results. Therefore, it becomes imperative for future research endeavors to delve into this aspect further and investigate how these factors might influence the generalizability of our findings across diverse linguistic contexts.

### v3

@@ 7.6 Auditory Decoding
With the recent advancements of self-supervised speech models and generative AI models, recent studies have largely targeted reconstructing speech/music from brain recordings (Défossez et al., 2023; Denk et al., 2023; Senda et al., 2024). As shown in Figure 20Défossez et al. (2023) proposed a CLIP-MEG pipeline to align MEG activity onto pretrained speech embeddings and generate speech from a stream of MEG signals. Unlike other methods which are experimented with on narrative speech, Denk et al. (2023) introduced a method for reconstructing music from fMRI brain activity, as shown in Figure 21. Specifically, they proposed a Brain2Music pipeline where the first step involves using fMRI data to predict MuLanmusic embeddings (Huang et al., 2022), which are then passed to MusicLM (Agostinelli et al., 2023), is conditioned to generate the music reconstruction, resembling the original music stimulus.
Figure 20: CLIP-MEG pipeline to align MEG activity onto pretrained speech embeddings. This Figure is adapted from (Défossez et al., 2023), and used with permission from the respective authors.
Figure 21: Brain2Music decoding pipeline to reconstruct Music from fMRI brain activity (Denk et al., 2023)This Figure is adapted from (Denk et al., 2023), and used with permission from the respective authors.
@@ 7.7.1 Image Reconstruction
Before the success of recent generative AI models, researchers have used deep-learning models and algorithms, including generative adversarial networks (GANs) and self-supervised learning models trained on a large number of naturalistic images (Du et al., 2020; Beliy et al., 2019; Fang et al., 2020; Gaziv et al., 2022; Lin et al., 2022). For instance, Beliy et al. (2019) designed a separable autoencoder that enables self-supervised learning in fMRI and images to increase training data.  Lin et al. (2022) introduced a framework, Mind Reader, that encoded fMRI signals into a pre-aligned vision-language latent space and used StyleGAN2 (Karras et al., 2020) for image generation. These methods generate more plausible and semantically meaningful images. Several other studies focused on reconstructing personal imagined experiences (Berezutskaya et al., 2020) or application-based decoding like using brain activity scanned during a picture-based mechanical engineering task to predict individuals’ physics/engineering exam results (Cetron et al., 2019) and reflecting whether current thoughts are detailed, correspond to the past or future, are verbal or in images (Smallwood & Schooler, 2015).
@@
With the recent success of CLIP and Diffusion models, deep generative models have been gaining attention to generate high-resolution images with high semantic fidelity (Takagi & Nishimoto, 2023; Chen et al., 2023; Scotti et al., 2024; Benchetrit et al., 2024; Song et al., 2023).  Takagi & Nishimoto (2023) proposed a method for image reconstruction from fMRI using Stable Diffusion (Rombach et al., 2022), as shown in Figure 22 (left). Their approach involves decoding brain activities to text descriptions and converting them to natural images using Stable Diffusion. Based on a similar philosophy, using a Stable Diffusion model as a generative prior and the pretrained fMRI features as conditions, Chen et al. (2023) reconstructed high-fidelity images with high semantic correspondence to the groundtruth stimuli, as shown in Figure 22 (right).  Scotti et al. (2024) proposed a MindEye that can map fMRI brain activity to any high dimensional multimodal latent space, like CLIP image space, enabling image reconstruction using generative models that accept embeddings from this latent space. Different from previous studies, BrainCLIP framework was introduced by Liu et al. (2023) to align fMRI patterns with different modalities (especially from visual and textual modalities) through cross-modal contrastive loss. All these studies have been limited to 2D visual representations. A recent work by Gao et al. (2023) aims to extend the scope of fMRI decoding to 3D representations. Specifically, they introduced Recon3DMind, a groundbreaking task focused on reconstructing 3D visuals from fMRI signals.
@@
Lastly, recent image reconstruction studies have focused on other non-invasive brain recordings such as MEG and EEG rather than fMRI signals.  Benchetrit et al. (2024) proposed a CLIP-MEG pipeline to align MEG activity with pretrained visual embeddings and generate images from a stream of MEG signals. Similarly, Song et al. (2023) proposed a CLIP-EEG pipeline to align these two modalities (image and EEG encoders to extract features from paired image stimuli and EEG responses) by constraining their similarity.
@@
Figure 22: Image reconstruction from fMRI using Stable Diffusion. Left: (Takagi & Nishimoto, 2023), Right: (Chen et al., 2024c)The Left Figure is adapted from (Takagi & Nishimoto, 2023) and the Right Figure is adapted from (Chen et al., 2024c). Adapted with permission from the respective authors.
@@ 7.7.2 Video Reconstruction
Unlike static natural images, human visual cortex can process a continuous, diverse flow of scenes, motions, and objects. To recover dynamic visual experience, the challenge lies in the nature of fMRI, which measures blood oxygenation level dependent (BOLD) signals and captures snapshots of brain activity every few seconds. Similar to image reconstruction works, Chen et al. (2024c) present MinD-Video, a two-module pipeline (i.e. CLIP module followed by latent stable diffusion) designed to bridge the gap between image and video brain decoding.
@@ 8 Conclusion, Limitations, and Future Trends
AI engineering. AI engineering generally refers to the design, development, and implementation of artificial intelligence models that simulate human cognitive functions (e.g., vision, language, memory) to enhance our understanding of how these processes are represented and processed in the brain. It also involves designing neural networks inspired by biological neural structures and principles, which can help uncover patterns associated with specific cognitive states or tasks.
The recent brain encoding studies most immediately fit in with the neuro-AI research direction that specifically investigates the relationship between representations in the brain and representations learned by powerful neural network models. This direction has gained recent traction, especially in the domain of language, vision, and speech processing, thanks to advancements in language models (Schrimpf et al., 2021; Goldstein et al., 2022), vision models (Schrimpf et al., 2020) and speech models (Tuckute et al., 2023; Oota et al., 2023c). Furthermore, several recent works most immediately contribute to this line of research by understanding the reasons for the observed similarity in more depth (Merlin & Toneva, 2022; Oota et al., 2024b; Kauf et al., 2024b; Sarch et al., 2024; Oota et al., 2024a). Overall, these studies provide valuable insights for selecting features, enhancing transfer learning, and aiding in the creation of AI architectures that are cognitively plausible.
@@
From humans to animal models. While this survey focuses on human neuroimaging, it is essential to recognize the significant role of animal models in neuroscience research. Studies involving rodents, primates, and even simpler organisms like C. elegans offer unique insights that complement human studies (Altevogt et al., 2012; Romanova & Sweedler, 2018). Animal models allow for genetic manipulation, invasive techniques, developmental studies, and evolutionary perspectives that are often impossible or unethical in human research. For instance, optogenetics, developed in animal models, enables precise control of neural circuits, providing causal insights into brain function (Romanova & Sweedler, 2018; Metwally et al., 2024).
In the context of encoding and decoding, many fundamental principles were first discovered in animal models. The concept of receptive fields in visual processing, for example, was initially described in cat visual cortex by Hubel & Wiesel (1962). Similarly, our understanding of place cells and grid cells in spatial navigation comes primarily from rodent studies but has implications for human spatial cognition (Moser et al., 2008). When interpreting results, especially in comparative studies, the species-specific nature of neural organization must be considered. However, it is crucial to note that findings from animal studies may not always directly translate to humans due to differences in brain structure and complexity (Kandel et al., 2000). By acknowledging this diversity in neuroscience research, we aim to provide a comprehensive understanding of neural information processing across species, while maintaining our focus on human neuroimaging studies in this review.
Ethical considerations and regulatory framework. It is crucial to understand the ethical landscape surrounding human brain research. The collection and use of neuroimaging data involves significant privacy and ethical considerations. Researchers must adhere to strict protocols to protect participants’ personal information, including rigorous anonymization processes to ensure that brain data cannot be linked back to individuals (Eklund et al., 2016). Many funding bodies and journals now mandate open-source data sharing to promote scientific transparency and reproducibility (Poldrack & Gorgolewski, 2014), but this must be balanced with privacy concerns. Before any study begins, it must be approved by Institutional Review Boards (IRBs) or Ethics Committees, which evaluate the potential risks and benefits to participants (Klitzman, 2013). These committees ensure that studies comply with national and international regulations, such as the General Data Protection Regulation (GDPR) in Europe or the Health Insurance Portability and Accountability Act (HIPAA) in the United States (Mourby et al., 2018). Informed consent is a cornerstone of ethical research, requiring that participants fully understand the nature of the study, its risks, and how their data will be used (Nijhawan et al., 2013). As neurotechnology advances, new ethical challenges emerge, such as the potential for brain data to reveal sensitive personal information beyond the scope of the original research question (Ienca & Andorno, 2017). The ongoing dialogue between researchers, ethicists, and policy makers is essential to navigate these complex issues and maintain public trust in neuroscience research.
Challenges and concerns in neuroscience community. In the neuroscience community, preprocessing brain data is a critical step in both functional Magnetic Resonance Imaging (fMRI) and Magnetoencephalography (MEG) data analysis. Effective preprocessing ensures that the data are clean, standardized, and suitable for subsequent analyses. However, both fMRI and MEG preprocessing pipelines face a variety of challenges and issues that can impact the quality and reliability of the results. Below is an overview of the common issues encountered in fMRI and MEG preprocessing:
  • Issues in fMRI Preprocessing Pipeline:
    1. 1.
      Motion Artifacts and Correction: Head movements during scanning can introduce significant artifacts, leading to spurious signals that mimic neural activity.
    2. 2.
      Physiological Noise: Fluctuations in heart rate and breathing can introduce noise into the fMRI signal.
    3. 3.
      Slice Timing Correction: Variations in the acquisition time of different slices within a single volume can lead to temporal misalignments.
    4. 4.
      Signal-to-Noise Ratio (SNR): Low SNR can obscure true neural signals, making it difficult to detect meaningful activations.
  • Issues in MEG Preprocessing Pipeline:
    1. 1.
      Environmental Noise and Interference: MEG is highly sensitive to external magnetic and electrical noise, which can contaminate the neural signals.
    2. 2.
      Source Reconstruction Complexities: Accurately reconstructing the sources of neural activity from sensor data is computationally intensive and prone to errors.
    3. 3.
      Integration with MRI Data: Combining MEG data with anatomical MRI data for accurate source localization and spatial referencing can be challenging.
Addressing these challenges requires continuous advancements in preprocessing frameworks such as fMRIPrep (Esteban et al., 2019) and MNE (Gramfort et al., 2013), the adoption of standardized preprocessing pipelines, and rigorous quality assurance protocols. By mitigating these preprocessing issues, researchers can improve the accuracy and reproducibility of code, data, and results in neuroimaging studies, thereby deepening our understanding of brain function and its underlying mechanisms. Ultimately, overcoming these preprocessing challenges is essential for advancing the development of cognitively plausible deep neural network (DNN) models and for enhancing practical applications in neuroscience.
Challenges in selecting evaluation metrics for brain-model alignment. Selecting the most appropriate metrics for assessing the alignment between computational models and brain data is an active area of debate in the neuroscience and machine learning communities (Williams et al., 2021; Soni et al., 2024). Each evaluation metric discussed has its own set of assumptions and limitations, which can lead to differing interpretations of model performance.
Recent studies have highlighted the complexities and subjective choices involved in metric selection. For instance, Williams et al. (2021) discuss the limitations of correlation-based metrics, such as PCC and $R^{2}$ score, suggesting that relying solely on them may provide an incomplete understanding of brain-model alignment. Similarly, debates have arisen over the use of distributional similarity measures like RSA, CKA, and NDS (Marques et al., 2021), which compare the statistical properties of neural and model representations but may not capture fine-grained functional correspondences. Recently, Soni et al. (2024) conducted a comprehensive evaluation of the aforementioned metrics across multiple brain datasets and models to examine the relationships between these measures and the extent to which they support similar conclusions. They found that not only do these measures cause inconsistencies in brain alignment within and across models, but these inconsistencies also undermine some of the conclusions made in the neuro-AI field. Overall, Soni et al. (2024) concluded that evaluation measures were often not chosen based on an analysis of the specific problem at hand but rather for compatibility with the wider field or due to mathematical elegance. However, considerable work may be needed to produce interpretable insights from distinct results obtained by using multiple measures. Further, it is essential to publish complete data and code for reproducibility and transparency. Any claims that rely solely on similarity scores should be revisited with a closer examination of the models.
@@ 8.1 Future Trends
Bridging the gap: enhancing deep neural network models for deeper insights into auditory, language and visual processing. While significant progress has been made in understanding text-based models, understanding the similarity in information processing between visual, speech, and multimodal models versus natural brain systems remains an open area. For instance, Oota et al. (2024a) demonstrate that speech-based language models lack brain relevant semantics in language regions. Therefore, enhancing speech-based language models to align more closely with text-based models could provide valuable insights into language and auditory processing, given that speech is the most ancient form of human language. This suggests a promising direction for future research, aiming to bridge the gap between artificial intelligence models and the complex, multifaceted processes of human cognition.
@@
Advancing multimodal decoding: the next leap in deep learning accuracy. Decoding actual multimodal stimuli has become increasingly feasible due to recent advancements in deep learning models dedicated to generation tasks (Rombach et al., 2022; Singer et al., 2022). However, there is still a significant need for further research to enhance the accuracy of these models. This involves not only refining the algorithms and architectures used but also improving the quality and diversity of the datasets on which these models are trained. Advancements in computational power, algorithmic efficiency, and innovative training methodologies are critical for pushing the boundaries of what is possible in multimodal decoding, aiming to achieve more precise, reliable, and nuanced interpretations of complex stimuli.
@@
Mapping the mind: the effects of brain damage on cognitive capabilities. We need a deeper understanding of the degree to which damage to different regions of the human brain could lead to the degradation of selective cognitive skills. This exploration requires detailed mapping of cognitive functions to specific brain areas, taking into account the brain’s complex network of connections. Studies should investigate not only the immediate effects of brain damage on cognitive skills but also the brain’s capacity for reorganization and compensation over time. Ultimately, the goal is to translate these research findings into practical applications, such as more effective cognitive rehabilitation techniques and assistive technologies to improve the quality of life for individuals with brain injuries.
@@
Towards human-like understanding in ANNs: integrating self-supervised learning and brain-inspired architectures. How can we train artificial neural networks in novel self-supervised ways such that they compose word meanings or comprehend images and speech like a human brain? Can we model the hierarchical and modular organization of the brain in neural network architectures? This involves creating networks that reflect the brain’s organization, from low-level feature detection to high-level semantic processing, allowing for the integration of information across different modalities. Moreover, how might we integrate dynamic learning strategies, such as curriculum learning, which progressively introduces more intricate tasks to the model? This method emulates how humans naturally progress from understanding straightforward to more complex ideas over time.
@@
Bridging the language gap in brain-NLP research: the need for multilingual exploration. An important part of brain-NLP research relies on brain recordings collected from individuals who speak English as their primary language. These studies also utilize experimental stimuli that are presented in the English language. As a result, the majority of neuro-AI research predominantly leverages language models and neural models that have been extensively trained on English text data and brain responses elicited by text or speech in English. However, it is essential to acknowledge the potential variability in study outcomes when extrapolated to languages other than English. For instance, a recent study by Chen et al. (2024b) focuses on recording brain responses while participants fluent in both English and Chinese read several hours of natural narratives in each language. Moreover, the availability of naturalistic stories such as Audiocite.net corpus in French (Felice et al., 2024), and The Little Prince audiobook in Cantonese (Momenian et al., 2024) provides suitable material for recording brain data in other languages. Therefore, it becomes imperative for future research endeavors to delve further into this aspect and investigate how these factors might influence the generalizability of our findings across diverse linguistic contexts.
@@
Neural population control: the next frontier. While significant progress has been made in encoding and decoding brain signals, a promising complementary direction involves controlling or stimulating neural activity to influence perception and behavior. This bidirectional brain-computer interface approach could facilitate not only reading but also writing information to the brain, with current methods ranging from optogenetics for precise control of neural circuits (Deisseroth, 2011) to electrical stimulation through neural prosthetics (Basumatary & Hazarika, 2020), and non-invasive techniques such as focused ultrasound and transcranial magnetic stimulation (Polanía et al., 2018). These approaches show particular promise in visual applications, such as restoring sight in blind individuals through direct stimulation of the visual cortex (Beauchamp et al., 2020) or enhancing visual perception through targeted neural modulation (Romei et al., 2010). Additionally, insights from deep-learning-based analyses suggest that targeted stimulation, such as presenting highly activating images, could not only control but also decode neural activity, using brain stimulation as a means to deepen our understanding of brain responses and their similarity to deep network activities (Bashivan et al., 2019; Ponce et al., 2019). Recent efforts using linguistic stimuli to drive language network activity (Tuckute et al., 2024) or use explanation-mediated language stimulus synthesis (Antonello et al., 2024a) show promise in controlling language ROIs. However, several technical challenges persist: the precise targeting of specific neural populations remains difficult, potential tissue damage from invasive methods raises concerns, and our incomplete understanding of neural codes complicates the determination of effective stimulation patterns. Despite these challenges, this direction naturally extends the decoding work discussed in this survey and could eventually lead to fully bidirectional brain-computer interfaces that can both interpret and influence neural activity in real-time (Saha et al., 2021). Such technology holds profound implications for treating neurological conditions, enhancing human capabilities, and advancing our understanding of brain function (Bashivan et al., 2019), provided that future research successfully addresses the needs for more precise and less invasive stimulation methods while carefully considering the ethical implications of direct neural control.